# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ VII Conclusions and Perspectives
We end this review by discussing open questions in this fast-developing field.
@@
Topological Phonons. Like electronic systems, topological phonons can give rise to phononic edge mode. It has been predicted that the double helicoidal surface phonon modes will emerge in transition metal monosilicides, such as FeSi and CoSi Zhang et al. (2018a); Miao et al. (2018); Jin et al. (2022), whereas the flat surface phonon modes can be realized in MoB2 Zhang et al. (2019b) and high-$T_{c}$ conventional superconductor MgB2 Li et al. (2020c). Experimental observation of these novel topological edge modes will be important for the understanding of topological bosonic excitations and related physical consequences. Surface sensitive probes, such as HR-EELS, will be a powerful technique to resolve this issue. IXS with grazing incident condition may also provide key insights.
@@
Furthermore, the coupling between topological phonon and topological electron remain largely unexplored. These couplings are expected to be important in Weyl semimetals, such as CoSi, where both Weyl phonon and Weyl fermions are crucial for transport properties. The coupling between Weyl electron and phonon can give rise to non-linear Hall effect He et al. (2019); Gao et al. (2020). The topological quasiparticle interactions can also be important for spontaneous symmetry-breaking orders, such as charge/spin density waves Luo et al. (2018); Miao et al. (2024); Yang et al. (2024a); Zhang et al. (2024b), quantum Hall effects Luo et al. (2018), and superconductivity Grissonnanche et al. (2020); Gao et al. (2023). The localized surface/edge states induced by nontrivial bulk band topology support peculiarly large density of states for both electrons and phonons, which may give rise to enhanced electron-phonon couplings at some interfaces. Di Miceli et al. (2022) has shown that topological phonons can play important roles in superconductivity. Investigations along this direction may help deepen our understanding about phonon-mediated superconductivity. Understanding of these topological quasiparticle interactions may also pave the way for practical applications, e.g., nonreciprocal transport and enhanced thermoelectric conversion efficiency Xu et al. (2014).
@@
Chiral Phonons. Previous experimental studies suggest that chiral phonons may couple to magnetic degree of freedom. However, quantitative understanding of the coupling remains unresolved. For instance, it has been known that chiral structure can induce spin polarized current Ray et al. (1999); Göhler et al. (2011); Naaman and Waldeck (2012); Naaman et al. (2019). A recent theoretical study showed that this chirality-induced spin selectivity arises from spin-orbital coupling, which polarizes electron spin rather than filters spin Wolf et al. (2022).
The coupling between chiral phonons and magnetism may also help to shed light in the unusual thermal transport behaviors in correlated and topological materials. As one of the main heat carriers, phonons are critical for thermal conductivity. For a long time, the phonon contribution to the thermal Hall conductivity, $\kappa_{xy}$, has been overlooked due to the small phonon magnetic moment and hence negligible responses under magnetic fields. Recently, giant $\kappa_{xy}$ was observed in many strongly correlated insulators, including the cuprate high-Tc superconductors Grissonnanche et al. (2020), Kitaev spin liquid candidate $\alpha$-RuCl3 Kasahara et al. (2018), quantum paraelectric SrTiO3 Li et al. (2020d). Empirically, THE in these insulators displays two ubiquitous features: (1) longitudinal and transverse thermal conductivity peak at the same temperature, and (2) $\kappa_{xy}/\kappa_{xx}\sim 10^{-3}$ around Bsimilar-$\sim$10 T. These observations suggest that the unusually large THE has a phonon origin. Phenomenologically, the transverse conductivity can be written as $\kappa_{xy}=\frac{1}{\nu}\sum_{\gamma}C^{\gamma}(v_{xy}^{\gamma}l^{\gamma}+l_{xy}^{\gamma}v^{\gamma})$, where $C$, $v$, and $l$ are the specific heat, phonon velocity, and mean-free path, respectively. $\gamma$ is the index of phonon modes, and $\nu$ is a dimension-dependent constant. Like electrons, the phonon THE has intrinsic contribution from phonon Berry curvature Qin et al. (2012); Zhang et al. (2010) and extrinsic contribution from skew scattering and side-jump Chen et al. (2020). However, it has been found that the intrinsic effect is usually orders of magnitude smaller than experimental observations. These experimental observations call for further theoretical study of the coupling between phonons and their embedded environments Chen et al. (2022); Li et al. (2023b); Zhang et al. (2019c); Grissonnanche et al. (2019).
@@
In summary, topological and chiral phonons represent a transformative frontier in condensed matter physics. These unconventional lattice excitations exhibit nontrivial topology and carry (pseudo-)angular momentum that are important to understand light-matter scattering, thermal transport, and emergent symmetry breaking and topological orders. These concepts expand the understanding of symmetry, topology, and transport in crystalline solids, with promising implications for both fundamental science and quantum technologies.
@@ VIII Acknowledgements
T.Z. acknowledges the support from the National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1400036) and the National Natural Science Foundation of China (Grant Nos. 12374165 and 12047503). Y.L. is supported by the Innovation Program for Quantum Science and Technology (Grant No. 2023ZD0300500), the National Natural Science Fundation of China (Grant No. 12404279), and Shanghai Pujiang Program (Grant No. 23PJ1413000). H.M. is supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. S.M. is supported by JSPS KAKENHI Grant Nos. JP22H00108 and JP24H02231.

### v2

@@ VII Conclusions and Perspectives
We end this review by summarizing main theoretical and experimental achievements and discussing open questions in topological phonons and circularly polarized phonons.
@@
Topological Phonons. The theory of topological phonons is derived from the topological band theory hasan2010colloquium ; Qin2011Nov ; Bansil2016Jun . Theoretical predictions motivated experimental studies to directly measure Dirac and Weyl phonons and chiral phononic wave functions in the bulk states. Like electronic systems, topological phonons can give rise to phononic edge mode. It has been predicted that the double helicoidal surface phonon modes will emerge in transition metal monosilicides, such as FeSi and CoSi zhang2018double ; miao2018observation ; jin2022chern , whereas the flat surface phonon modes can be realized in MoB2 zhang2019phononic and high-$T_{c}$ conventional superconductor MgB2 li2020phononic . Experimental observation of these novel topological edge modes will be important for the understanding of topological bosonic excitations and related physical consequences. Surface sensitive probes, such as HR-MEELS, will be a powerful technique to resolve this issue. RIXS and IXS with grazing incident condition may also provide key insights. With new experimental capabilities enabling selective excitation and detection of phonon modes, further topological phonon-driven phenomena in solids are likely to emerge. These advances also motivate applications in metamaterials, exemplified by topological phonon polaritons Guddala2021Oct and robust phononic devices such as disorder-immune delay lines and one-way waveguides Zhang2018Mar ; Cha2018Dec ; Xi2025Jun .
@@
An largely unexplored open question is the coupling between topological phonons and topological electrons. These couplings are expected to be important in Weyl semimetals, such as CoSi, where both Weyl phonon and Weyl fermions are crucial for transport properties. These coupling can give rise to non-linear Hall effect he2019nonlinear ; gao2020second . The topological quasiparticle-interactions can also be important for spontaneous symmetry-breaking orders, such as charge/spin density waves luo2018topological ; Miao2024Spontaneous ; yang2024incommensurate ; zhang2024understanding , quantum Hall effects luo2018topological , and superconductivity grissonnanche2020chiral ; gao2023chiral . The localized surface/edge states induced by nontrivial bulk band topology can help to enhance electron-phonon couplings and stabilize superconductivity at the interfaces DiMiceli2022Aug . Understanding of these topological quasiparticle-interactions may also pave the way for practical applications, e.g., nonreciprocal transport and enhanced thermoelectric conversion efficiency Xu2014Jun .
Another interesting open question is the quantum geometry of phonons. Since the imaginary part of quantum geometry is the Berry curvature, phonons are expected to have non-trivial quantum metrics, the real part of quantum geometry provost1980riemannian ; berry1989quantum ; berry1984quantal ; kim2025direct ; yu2020experimental ; tan2019experimental ; gianfrate2020measurement ; kang2025measurements ; bleu2018effective ; neupert2013measuring ; albert2016geometry ; yu2024non . Recently, it has been shown that under the Gaussian approximation, the lower bound of electron-phonon coupling is determined by the quantum geometry of the electronic structure yu2024non . The role of quantum geometry of phonons in the electron-phonon coupling is yet to be understood.
Circularly polarized Phonons. Unlike topological phonons, circularly polarized phonons have much longer history. The main conceptual progress in this field is the development of PAM that can be directly measured using polarization controlled optical probes zhu2018observation ; ishito2023chiral ; zhang2023weyl ; zhang2025thechirality ; che2025magnetic ; yang2025inherent ; watanabe2025symmetry . As we described in Section V, the PAM is different from AM of phonons. Unlike the PAM that can be determined directly through Raman and IR probes, the phonon AM is measured until very recently using a mechanical torque setup zhang2024observation .
@@
While AM can be transferred between phonons and electronic excitations through quasiparticle-interactions, the magnitudes of these scattering channels remain unresolved, which may complicate the interpretations of experimental data. For instance, it has been known that chiral structure can induce spin polarized current ray1999asymmetric ; gohler2011spin ; naaman2012chiral ; naaman2019chiral . A recent theoretical study showed that this chirality-induced spin selectivity arises from spin-orbital coupling, which polarizes electron spin rather than filters spin through AM transfer from chiral phonon wolf2022unusual .
[…]
Phonons with Both Topological Chirality and Rotational Chirality. Breaking both $\mathcal{P}$ and $\mathcal{T}$ symmetries is the essential condition for realizing nonzero Chern number and angular momentum zhang2018double ; zhang2025weyl . Under open boundary conditions or at the interface between different domains, topological surface phonons, that are robust against impurities and structural imperfections, can emerge. For example, studies of thermal transport in chiral $\alpha$-quartz reveal that phonons can carry chirality-dependent angular momentum ohe2024chirality , which can be extracted as a spin signal. This effect suggests that topological surface phonons may also exhibit chirality-dependent angular momentum, potentially contributing to robust spin and heat transport. Furthermore, introducing magnetic or other $\mathcal{T}$-breaking fields, can induce topological and rotational chirality even centrosymmetric materials, leading to nonreciprocal phonon transport hirokane2020nonreciprocal . The topological protection ensures low-dissipative surface phonons, while the rotational chirality enables angular momentum transfer. Consequently, the longitudinal thermal conductivity becomes dependent on whether the thermal current is parallel or antiparallel to the vector product of electric polarization and magnetization, paving the way for phonon-based thermal diodes with controllable and scalable functionality.
@@
In summary, topological and circularly polarzied phonons represent a transformative frontier in condensed matter physics. These unconventional lattice excitations exhibit nontrivial topology and carry nonzero AM and PAM that are important to understand light-matter scattering, thermal transport, and emergent symmetry breaking orders and topological orders. These concepts expand the understanding of symmetry, topology, and transport in crystalline materials, with promising implications for both fundamental science and quantum technologies.
@@ VIII Acknowledgements
T.Z. acknowledges the support from the National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200) and the National Natural Science Foundation of China (Grant Nos. 12374165 and 12047503). Y.L. is supported by the Innovation Program for Quantum Science and Technology (Grant No. 2023ZD0300500), the National Natural Science Fundation of China (Grant No. 12404279), and Shanghai Pujiang Program (Grant No. 23PJ1413000). H.M. is supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. S.M. is supported by JSPS KAKENHI Grant Nos. JP22H00108 and JP24H02231.